# Emergence of non-Markovian Decoherent Histories in Integrable Environment: A “Tape Recorder” Model for Local Quantum Observables

### v2

@@ VI Unraveling of dynamics
A particularly appealing consequence of the decoherent histories framework is that the quantum evolution can be interpreted as a classical stochastic process on the space of decoherent histories. In this sense, classical reality is encoded in the ensemble of emergent decoherent histories.
@@
Figure 8 was obtained in this way, by averaging over 100 randomly generated histories.
Figure 8: Time evolution of $\langle\sigma_{z}\rangle$. The irreversibly decoupled modes are traced out using the quantum jump Monte Carlo simulation (Sec. V), providing an efficient unraveling of the full wave function into a sampled ensemble of decoherent histories.
@@ VII Discussion
We have shown that the tape-recorder coarse-graining provides a physically grounded way to construct decoherent histories even in integrable, non-Markovian environments. Although our explicit analysis was limited to a spin-boson model, the underlying mechanism is quite general, and we expect that it can be extended to other types of environments, including chaotic case.
Current research on out-of-equilibrium many-body quantum dynamics focuses on questions such as how quenched systems approach equilibrium; how the correlations and quantum information propagate throughout the system; how such factors as integrability, disorder affect the quench dynamics. Concepts such as operator growth, entanglement spreading, and out-of-time-order correlators (OTOCs) play a central role in these studies. Our approach complements it and provides a framework for explicitly constructing multi-time decoherent histories, in which the system evolves along a branching structure of mutually exclusive alternatives.
In this construction, the key carriers of information are the irreversibly decoupled environmental modes, which act as stable records of the system’s past. Each history corresponds to an effective sequence of quantum jumps, so that the full many-body dynamics can be interpreted as an unraveling into a branching ensemble of trajectories. This also enables efficient Monte Carlo sampling of histories, as illustrated by the agreement between observables computed via exact dynamics and those reconstructed from history unravelings (Fig. 8).
Beyond the explicit construction of histories, the tape-recorder coarse-graining naturally opens new perspectives. In particular, in earlier work we defined the entropy of decoherent histories and argued that it can serve as a diagnostic of quantum chaos PolAref:24 . Establishing a systematic connection to non-equilibrium thermodynamics remains an open question for future study.
@@ VIII Conclusions
We introduced a mechanism for the emergence of decoherent histories in integrable, non-Markovian environments, based on viewing the environment as a tape recorder that sequentially and irreversibly stores information about the system. The approach provides a controllable way to construct consistent multi-time histories without assuming non-integrability or Markovianity.
@@ Appendix A Infidelity bound and justification of the Lieb-Robinson metric
$\hat{a}_{0}^{\dagger}\left(t\right)=\hat{a}_{\perp}^{\dagger}\left(t\right)+\chi\left(t\right)\hat{\kappa}^{\dagger}{}^{\rm edge}$, (48)
@@
$\sqrt{F(\kappa,t)}=\left|\langle\Psi_{\perp}(t)|\Psi(t)\rangle\right|$ (51)
@@
$I\left(\kappa,t\right)=1-\sqrt{F\left(\kappa,t\right)}$ (52)
provides a measure of the statistical significance of the mode $\kappa$ over the interval $[0,t]$.
@@
$I\left(\kappa,t\right)=I_{0}\left(t\right)+I_{2}\left(\kappa,t\right)+\ldots+I_{2p}\left(\kappa,t\right)+\ldots.$ (53)
Only even-order terms appear in this expansion because the vacuum expectation values of odd powers of $\hat{\kappa},\hat{\kappa}^{\dagger}$ vanish in eq. (51). Therefore, each $\hat{\kappa}^{\dagger}$ should be annihilated by some $\hat{\kappa}$, which contributes factor $\chi^{*}\left(\tau^{\prime}\right)\chi\left(\tau\right)$. At order zero we have $\Psi\left(t\right)=\Psi_{\perp}\left(t\right)$, therefore $I_{0}\left(t\right)=0$. The term $I_{2p}\left(\kappa,t\right)$ can be written as:
$I_{2p}\left(\kappa,t\right)=\left[\frac{1}{p!}\right]^{2}\left\langle\Psi_{\perp}\left(t\right)\right|\intop_{0}^{t}d\tau_{1}\ldots d\tau_{p}d\tau_{1}^{\prime}\ldots d\tau_{p}^{\prime}\mathcal{T}\prod_{i=1}^{p}\left[g\chi^{*}\left(\tau_{i}\right)\hat{V}_{s}^{\dagger}\hat{\kappa}\right]\prod_{j=1}^{p}\left[g\chi\left(\tau_{j}^{\prime}\right)\hat{V}_{s}\hat{\kappa}^{\dagger}\right]\exp\left[-i\intop_{0}^{t}d\tau\hat{H}_{\perp}\left(\tau\right)\right]\left|\Psi\left(0\right)\right\rangle.$ (54)
@@
$\left|I_{2p}\left(\kappa,t\right)\right|\leq\left[\frac{1}{p!}\right]^{2}p!\left[g\left\|\hat{V}_{s}\right\|\right]^{2p}\left\|\chi\right\|_{L^{2}}^{2p},$ (55)
@@
$\left\|\chi\right\|_{L^{2}}^{2}=\intop_{0}^{t}d\tau\left|\chi\left(\tau\right)\right|^{2}=\intop_{0}^{t}d\tau\left|\left\langle\kappa\left|\alpha\left(t\right)\right.\right\rangle\right|^{2}=\left\langle\kappa\right|\hat{\rho}_{+}\left(t\right)\left|\kappa\right\rangle,$ (56)
@@
$I\left(\kappa,t\right)\lesssim\left[g\left\|\hat{V}_{s}\right\|\right]^{2}a_{\rm cut},$ (59)
@@
$I\left(\kappa,t\right)\lesssim\left[g\left\|\hat{V}_{s}\right\|\right]^{2}\mathcal{I}_{1}^{+}\left(\infty\right)e^{-\gamma\delta}$ (61)
@@ Appendix B Relevant subspace for local quench on a finite time interval
Suppose we consider the evolution eq. (15-16) on a finite time interval $\left[0,T\right]$. At the final time $T$ we find the light cone interior normal modes $\left|\widetilde{\kappa}_{1}\right\rangle\ldots|\widetilde{\kappa}_{m\left(T\right)}\rangle$ as described in Sec. IV.1. In fact, additional normal modes: the eigenstates $|\widetilde{\kappa}_{m\left(T\right)+p}\rangle$ for $p>0$ fall below the significance threshold, violating condition (31). Owing to the light-cone consistency property (25), these modes remain below the threshold for all times $t\in\left[0,T\right]$. Consequently, by choosing a sufficiently small threshold $a_{\rm cut}$, we can safely discard these modes from the Hamiltonian (13) with only negligible distortion of computed observables $\propto O\left(a_{\rm cut}\right)$ throughout the interval $\left[0,T\right]$.
[…]
@@
$\hat{a}_{0}^{\dagger}\left(t\right)=\sum_{p=1}^{\infty}\alpha_{p}\left(t\right)\sum_{k=1}^{\infty}\left[\widetilde{\kappa}_{k}^{*}\right]_{p}\hat{\widetilde{\kappa}}_{k}^{\dagger}=\sum_{k=1}^{\infty}\left\langle\widetilde{\kappa}_{k}\left|\alpha\left(t\right)\right.\right\rangle\hat{\widetilde{\kappa}}_{k}^{\dagger}=\sum_{k=1}^{m\left(T\right)}\left\langle\widetilde{\kappa}_{k}\left|\alpha\left(t\right)\right.\right\rangle\hat{\widetilde{\kappa}}_{k}^{\dagger}+\sum_{k=m\left(T\right)+1}^{\infty}\left\langle\widetilde{\kappa}_{k}\left|\alpha\left(t\right)\right.\right\rangle\hat{\widetilde{\kappa}}_{k}^{\dagger}$ (63)

### v3

@@ VI Unraveling of dynamics
A particularly appealing consequence of the decoherent histories framework is that the quantum evolution admits an approximate stochastic unraveling in the space of decoherent histories. In this sense, classical reality is encoded in the ensemble of emergent decoherent histories.
@@
Figure 9 was obtained in this way, by averaging over 300 randomly generated histories for different significant threshold $r_{\rm cut}$, defining the number of relevant modes.
@@ VII Discussion
We have shown that the tape-recorder coarse-graining provides a physically grounded way to construct decoherent histories in integrable, non-Markovian environments. Although our explicit analysis was performed for a spin-boson model, the underlying mechanism is quite general and we expect that it can be extended to a broad class of open quantum systems, including those with chaotic environments.
Current research on out-of-equilibrium many-body quantum dynamics focuses on questions such as how quenched systems approach equilibrium; how the correlations and quantum information propagate throughout the system; how factors such as integrability, disorder affect the quench dynamics Eisert2015 . Concepts such as operator growth, entanglement spreading, and out-of-time-order correlators (OTOCs) Xu:2022 ; Hashimoto:2017 play a central role in these studies. Our approach complements it and provides a framework for explicitly constructing multi-time decoherent histories, in which the system evolves along a branching structure of mutually exclusive alternatives.
The key carriers of information in this construction are the irreversibly decoupled environmental modes, which act as stable records of the system’s past. Each history corresponds to an effective sequence of quantum jumps, so that the full many-body dynamics can be interpreted as an unraveling into a branching ensemble of trajectories. This enables efficient Monte Carlo sampling of histories, as illustrated by the agreement between observables computed via exact dynamics and those reconstructed from history unravelings (Fig. 9).
The significance threshold $r_{\rm cut}$, which determines the coarse-graining of the environment, is not an artificial parameter, it plays a role analogous to the noise floor of the recording head in the tape recorder picture. Its physically motivated and practically optimal choice is the experimentally achievable infidelity. Just as a tape recorder cannot meaningfully distinguish a signal below its noise floor, quantum superpositions with residual coherences below the infidelity are operationally indistinguishable from classical mixtures.
With this choice, the off-diagonal elements of the decoherence functional — the interference terms between distinct histories — are suppressed to order $O(\sqrt{r_{\rm cut}})=O(\sqrt{I})$, making the histories decoherent at the level of operational distinguishability. As a result, the construction is uniform in time and does not exhibit a breakdown time within the considered dynamics.
Beyond the explicit construction of histories, the tape-recorder coarse-graining naturally opens new perspectives. In particular, in earlier work we defined the entropy of decoherent histories and argued that it can serve as a diagnostic tool of quantum chaos PolAref:24 . Establishing a systematic connection to non-equilibrium thermodynamics remains an open question for future study.
@@ VIII Conclusions
We introduced a mechanism for the emergence of decoherent histories in integrable, non-Markovian environments, based on viewing the environment as a tape recorder that sequentially and irreversibly stores information about the system’s past. This provides a controllable and explicit way to construct consistent multi-time histories without assuming non-integrability or Markovianity. Decoherence is controlled by a physically meaningful threshold $r_{\rm cut}$ set by the experimentally accessible infidelity, and the resulting suppression of interference remains stable in time, with residual interference bounded by the infidelity. This makes the approach practically applicable beyond the short-time regime.
@@ Appendix A Infidelity bound and justification of the Lieb-Robinson metric
$\hat{a}_{0}^{\dagger}\left(t\right)=\hat{a}_{\perp}^{\dagger}\left(t\right)+\eta\left(t\right)\hat{\kappa}^{\dagger}{}^{\rm edge}$, (56)
@@
$\sqrt{F(t)}=\left|\langle\Psi_{\perp}(t)|\Psi(t)\rangle\right|$ (59)
@@
$I\left(t\right)=1-\sqrt{F\left(t\right)}$ (60)
provides a measure of the statistical significance of the mode $\kappa^{\rm edge}$ over the interval $[0,t]$. For brevity, we denote the edge mode $\kappa^{\rm edge}$ simply by $\kappa$.
@@
$I\left(\kappa,t\right)=I_{0}\left(\kappa,t\right)+I_{2}\left(\kappa,t\right)+\ldots+I_{2p}\left(\kappa,t\right)+\ldots.$ (61)
Only even-order terms appear in this expansion because the vacuum expectation values of odd powers of $\hat{\kappa},\hat{\kappa}^{\dagger}$ vanish in eq. (59). Therefore, each $\hat{\kappa}^{\dagger}$ should be annihilated by some $\hat{\kappa}$, which contributes factor $\eta^{*}\left(\tau^{\prime}\right)\eta\left(\tau\right)$. At order zero we have $\Psi\left(t\right)=\Psi_{\perp}\left(t\right)$, therefore $I_{0}\left(\kappa,t\right)=0$. The term $I_{2p}\left(\kappa,t\right)$ can be written as:
$I_{2p}\left(\kappa,t\right)=\left[\frac{1}{p!}\right]^{2}\left\langle\Psi_{\perp}\left(t\right)\right|\intop_{0}^{t}d\tau_{1}\ldots d\tau_{p}d\tau_{1}^{\prime}\ldots d\tau_{p}^{\prime}\mathcal{T}\prod_{i=1}^{p}\left[g\eta^{*}\left(\tau_{i}\right)\hat{V}_{s}^{\dagger}\hat{\kappa}\right]\prod_{j=1}^{p}\left[g\eta\left(\tau_{j}^{\prime}\right)\hat{V}_{s}\hat{\kappa}^{\dagger}\right]\exp\left[-i\intop_{0}^{t}d\tau\hat{H}_{\perp}\left(\tau\right)\right]\left|\Psi\left(0\right)\right\rangle.$ (62)
@@
$\left|I_{2p}\left(\kappa,t\right)\right|\leq\left[\frac{1}{p!}\right]^{2}p!\left[g\left\|\hat{V}_{s}\right\|\right]^{2p}\left\|\eta\right\|_{L^{2}}^{2p},$ (63)
@@
$\left\|\eta\right\|_{L^{2}}^{2}=\intop_{0}^{t}d\tau\left|\eta\left(\tau\right)\right|^{2}=\intop_{0}^{t}d\tau\left|\left\langle\kappa\left|\alpha\left(\tau\right)\right.\right\rangle\right|^{2}=\left\langle\kappa\right|\hat{\rho}_{+}\left(t\right)\left|\kappa\right\rangle,$ (64)
@@
$I\left(t\right)\lesssim\left[g\left\|\hat{V}_{s}\right\|\right]^{2}a_{\rm cut},$ (67)
@@
$I\left(t\right)\lesssim\left[g\left\|\hat{V}_{s}\right\|\right]^{2}\mathcal{I}_{1}^{+}\left(\infty\right)e^{-\gamma\delta}$ (69)
@@ Appendix B Relevant subspace for local quench on a finite time interval
Suppose we consider the evolution eq. (15-16) on a finite time interval $\left[0,T\right]$. At the final time $T$ we find the light cone interior normal modes $\left|\widetilde{\kappa}_{1}\right\rangle\ldots|\widetilde{\kappa}_{m\left(T\right)}\rangle$ as described in Sec. IV.1. In fact, additional normal modes: the eigenstates $|\widetilde{\kappa}_{m\left(T\right)+p}\rangle$ for $p>0$ fall below the significance threshold, violating condition (35). Owing to the light-cone consistency property (28), these modes remain below the threshold for all times $t\in\left[0,T\right]$. Consequently, by choosing a sufficiently small threshold $a_{\rm cut}$, we can safely discard these modes from the Hamiltonian (13) with only negligible distortion of computed observables $\propto O\left(\sqrt{a_{\rm cut}}\right)$ throughout the interval $\left[0,T\right]$.
@@
$\hat{a}_{0}^{\dagger}\left(t\right)=\sum_{p=0}^{\infty}\alpha_{p}\left(t\right)\sum_{k=1}^{\infty}\left[\widetilde{\kappa}_{k}^{*}\right]_{p}\hat{\widetilde{\kappa}}_{k}^{\dagger}=\sum_{k=1}^{\infty}\left\langle\widetilde{\kappa}_{k}\left|\alpha\left(t\right)\right.\right\rangle\hat{\widetilde{\kappa}}_{k}^{\dagger}=\sum_{k=1}^{m\left(T\right)}\left\langle\widetilde{\kappa}_{k}\left|\alpha\left(t\right)\right.\right\rangle\hat{\widetilde{\kappa}}_{k}^{\dagger}+\sum_{k=m\left(T\right)+1}^{\infty}\left\langle\widetilde{\kappa}_{k}\left|\alpha\left(t\right)\right.\right\rangle\hat{\widetilde{\kappa}}_{k}^{\dagger}$ (71)